STABILITY OF HO_{3} RADICAL VIA INFRARED ACTION SPECTROSCOPY

\underline{TIMOTHY D. SECHLER}, CRAIG MURRAY, ERIKA L. DERRO and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

The existence of hydrogen trioxy radicals (HO_{3}) has several important implications in atmospheric chemistry, as it may act as an intermediate in key chemical reactions and relaxation processes in the atmosphere. In this work, HO_{3} was produced from the association of O2 and photolytically generated OH radicals in the collisional region of a supersonic free-jet expansion. Excitation of the OH stretching vibration of HO_{3} via tunable infrared radiation causes dissociation to OH + O_{2}, and the OH fragments are probed by laser-induced fluorescence. The IR action spectrum of HO_{3} was recorded for the fundamental and first overtone OH stretches of HO_{3} as well as its deuterated analog. Rotationally structured and broadened features were observed and attributed to the trans and cis conformers of HO_{3}. The product state distribution of the OH fragment was examined to determine the highest energetically open product channel, which sets an upper limit of 6.12 kcal mol^{−1} for the HO-O_{2} binding energy. Calculation of the equilibrium constant for the reaction OH + O_{2} + M ⇀ ↽ HO_{3} + M (with collision partner M) suggests that HO_{3} exists in significant abundance in the vicinity of the tropopause, where it may act as a temporary sink for OH.